OpenStudy (anonymous):

Griselda budgets forty-five dollars per month before sales tax for clothes. How much will she spend on clothes each month if she is charged seven percent sales tax?

undeadknight26 (undeadknight26):

hmmm 45/.7

undeadknight26 (undeadknight26):

I believe it is that...

undeadknight26 (undeadknight26):

cuz it is definitely not multiplication, addition or subtraction!

OpenStudy (anonymous):

thirty-one dollars and fifty cents forty-one dollars and eighty-five cents forty-eight dollars and fifteen cents fifty-two dollars these are the answers

undeadknight26 (undeadknight26):

i think its the first one then...

undeadknight26 (undeadknight26):

guess it was multiplication...but how does sales tax decrease the amount of money you pay for something???

OpenStudy (jdoe0001):

$45 flat, she has to include that whatever her purchase will be, it will have a tax of 7% so she would have to include $45 as the amount she can use to purchase something, and that amount has to include the 7% taxation already so her figure to spend before tax will be less than $45 or $45-(7% of $45)

OpenStudy (jdoe0001):

hmmm.... well.... wait... rather..

OpenStudy (jdoe0001):

her figure to spend will be, SOME AMOUNT that will be less than $45 and will include the 7% tax

OpenStudy (jdoe0001):

let's say she's going to spend "x" amount an "x" amount plus 7% of "x" will be 45 so x + (7% of x) = 45 meaning \(\bf x + 0.07x = 45\)

OpenStudy (jdoe0001):

\(\bf \textit{n% of m} = \cfrac{n}{100}\cdot m\qquad thus\quad \textit{7% of "x"}=\cfrac{7}{100}\cdot x\implies 0.07x\)

OpenStudy (anonymous):

I am lost

OpenStudy (jdoe0001):

if she spends say $45 at the store... and they charge her 7% tax.... so she would have to pay..... 45 + tax => 45 + (7% of 45) => 45 + 0.07*45 = 48.15 so she will be over her budget by $3.15 so whatever she will spend at the store, cannot be $45 in price BEFORE TAX, it has to be less

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

so once the tax kicks in, it rounds up to $45, which is what she has

OpenStudy (jdoe0001):

yw